2. A ball is thrown at an angle of 45.0ยฐ from the horizontal with an initial velocity of 5.0 ๐‘š/๐‘ .

a. At what time will the ball reach its maximum height?

b. What will be the maximum height of the ballโ€™s trajectory?

c. How far will the ball travel horizontally when it reaches the maximum height of its flight?


1
Expert's answer
2022-02-28T10:38:36-0500

a. At what time will the ball reach its maximum height?

"t_1=\frac{v_0\sin\theta}{g}=\frac{5.0\sin45^\circ}{9.8}=0.36\:\rm s"

b. What will be the maximum height of the ballโ€™s trajectory?

"h_{\max}=\frac{v_0^2\sin^2\theta}{2g}=\frac{5.0^2\sin^245^\circ}{2*9.8}=0.63\:\rm m"

c. How far will the ball travel horizontally when it reaches the maximum height of its flight?

"l=v_0\cos\theta*t_1=5.0\cos45^\circ*0.36=1.27\:\rm m"